A density measurement was conducted on a sample of copper, which had a theoretical density of 8.91 g/ml. The copper had a mass of 62.20 g. and was a cylinder, which had a length of 3.4 cm. and a diameter of 1.6 cm.

Show the calculation of the volume of the copper cylinder.

Show the calculation of the density of the copper.

Show the calculation of the percent error based on the theoretical density of 8.91 g/ml.
1
Expert's answer
2017-03-01T09:00:49-0500
Answer
V=S*h
S=¶r²;r=0.8sm
S=2.0096sm²
V=2.0096*3.4=6.83sm³
₰=m/V
₰=62.20g/6.83sm³=9.11g/sm³
%error=|8.91-9.11|*100%/8.91=2.24%